Joan has a solution that is 18% acid and a second solution that is 43% acid. How many liters of each should she mixed together to get an 80L solution that is 26.5% acid?
Answer by Alan3354(69443)

Joan has a solution that is 18% acid and a second solution that is 43% acid. How many liters of each should she mixed together to get an 80L solution that is 26.5% acid?
---------------------
e = amount of 18%
f = amount of 43%
---
e + f = 80 ----- total solution
18e + 43f = 80*26.5 ---- total acid
------
Can you do the rest?